OpenStudy (anonymous):

your teacher is giving you a test worth 100 points containing 40 questions. there are two point questions and four point questions on the test. how many of each type of question are on the test. use a system of linear equations to solve

OpenStudy (anonymous):

Let 2 mark questions be x Let 4 mark questions be y

OpenStudy (anonymous):

Thus you get these two equations X+y=100 2x+4y=100

OpenStudy (anonymous):

ok I got that much

OpenStudy (anonymous):

Solve them simultaneously to get the x and y values

OpenStudy (anonymous):

Magic

OpenStudy (anonymous):

wait what happened to 40 and how do you solve simultaneously

OpenStudy (anonymous):

You got the question x+y=40

OpenStudy (anonymous):

That means the no. Of x and y question are 40 according to the data

OpenStudy (anonymous):

yeah you wrote x +y = 100

OpenStudy (anonymous):

Lolz

OpenStudy (anonymous):

My bad

OpenStudy (anonymous):

alright no its fine im just making sure. so then how would you solve them simultaneously?

OpenStudy (anonymous):

Do you know how to do simultaneous question?

OpenStudy (anonymous):

Wait nevermind ive got it! its x=30 and y=10. Nevermind were good. thank you

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

And good luck